Corporate Finance Class:

The Branding Iron Company sells its irons for $67 apiece wholesale. Production cost is $57 per iron. There is a 18% chance that wholesaler Q will go bankrupt within the next year. Q orders 1,000 irons and asks for eight months’ credit. Assume that the discount rate is 12% per year, there is no chance of a repeat order, and Q will pay either in full or not at all.

a. Calculate the NPV of the order. (A negative answer should be indicated by minus sign. Do not round intermediate calculations. Round your answer to 2 decimal places.)

NPV= $
Answer #1

Initial Outlay = Production Cost = Cost per Iron*Quantity = 57*1000 = $57000

Probability of going bankrupt within 8 months = Probability of going bankrupt within 1 year*2/3 = 18*2/3 = 12%

Expected Revenue = Revenue*Probability = (Revenue per unit*Quantity)*(1-Probability of going bankrupt in 8 months) = (67*1000)*(1-12%) = 67000*78% = $52260

PV of Expected Revenue = Expected Revenue/[1+(Interest Rate*2/3)] = 52260/[1+(0.12*2/3)] = 52260/1.08 = $48388.89

NPV = PV of Expected Revenue - Initial Outlay = 48388.89-57000 = $-8611.11